Question 3: Computations for fixed and variable costs At current sales volume of 100 units, fixed costs (FC) are $4 per unit and variable costs (VC) are $8 per unit. a) Compute total fixed costs at current sales volume. total FC400 b) Suppose that sales volume increases to 125 units. At this new volume, total FC FC per uni VC per unit = total VC Enter a number c) Write down the total cost equation: * volume (e.g., if TC 500+2*volume, enter 500 in the first box and 2 in the second box) Predict total costs at sales volume of 80 units: Predict total costs at sales volume of 120 units: d) When sales volume increases by 1 unit, total costs increase by unit VC $8 increase by unit FC $4 increase by unit cost $12 not enough information Submit Answer Save Progress Practice Another Version
Answer #1

a)Total Fixed Costs at current sales volume = 100*$4 = $400

b)Fixed Costs do not change with the change in volume of production.

Hence, If the sales volume increases to 125 units, Total FC = $400

FC per Unit = $400/125 = $3.2

Variable costs changes with the change in volume, VC per Unit = $8

Total VC = 125*$8 = $1,000

c) Total Cost Equation = Fixed cost + VC per unit*Volume

= $400 + $8*Volume

Predicted Total Cost at volume of 80 units = 400 + 8*80 = $1,040

Predicted Total Cost at volume of 120 units = 400 + 8*120 = $1,360

d)When Sales volume increases by 1 unit, Total Costs

Increase by Unit Variable Cost = $8